EXAMPLE - SQUARE ROOT Of -10, 6 SQRT OF 10
ANSWER - X^2 - 5 SQRT OF 10x - 60 = 0
(number 5 is on the outside of the SQRT. Maybe its a multiple? Wasn't sure how to explain.)
Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in the form, ax^2 + bx + c = 0
QUESTION: SQRT of -7, 5 SQRT 7
Thank you so much!!!

Answer by stanbon(75887)

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in the form, ax^2 + bx + c = 0
QUESTION: SQRT of -7, 5 SQRT 7
------
(x-sqrt(-7))(x-5sqrt(7)) = 0
----
(x-isqrt(7))(x-5sqrt(7)) = 0
---
x^2 -(5sqrt7+isqrt(7))x +7i = 0
---
x^2 -[sqrt(7)(5+i)]x + 7i = 0
===============================
Cheers,
Stan H.
===========